a tank can be filled by three pipes separaely in 20, 30, and 60 minutes respectively. In how many minutes can it be filled by the three pipes working together?
Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

a tank can be filled by three pipes separaely in 20, 30, and 60 minutes respectively. In how many minutes can it be filled by the three pipes working together?
.
Let x = time (minutes) it takes for all three
then
x(1/20 + 1/30 + 1/60) = 1
multiplying both sides by 60:
x(3 + 2 + 1) = 60
x(6) = 60
x = 60/6
x = 10 minutes

Answer by solver91311(24713)



If A can do a job in x time periods, then A can do of the job in 1 time period. Likewise, if B can do the same job in y time periods, then B can do of the job in 1 time period. For this problem you will need C doing of the job in 1 time period

So, working together, they can do



of the job in 1 time period.

Therefore, they can do the whole job in:



time periods.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism